If a car was to travel a distance of 180km at a speed of 60km/h, how long would this take? I know the formula is Time = Distance divided by Speed, but i did this and i got 3 minutes...surely this can't be right?...
Found 2 solutions by Alan3354, sachi:
Answer by Alan3354(69443)

If a car was to travel a distance of 180km at a speed of 60km/h, how long would this take? I know the formula is Time = Distance divided by Speed, but i did this and i got 3 minutes...surely this can't be right?...
-------------
The speed is km/hr, not km/minute
--> 3 hours.
==================
t = d/r = 180 km/60 km/hr
t = 3/(1/hr) = 3 hrs
Include the units in your calculations, it will improve the chances of getting it right.

Answer by sachi(548)

you are right but the time is hr not in mins
so the ans is 3 hrs
thanx